Journal entry worksheet 1 Record the issuance of 240,000 shares of common stock with a $1 par value for a price of $27 per shJournal entry worksheet 2 Record the purchase of 5,400 previously issued shares of common stock for a price of $28 per share.
Answer #1
Journal entries
Transaction Accounts Debit Credit Calculation
a Cash    6,480,000 =240,000*27
Common Stock       240,000 =240,000*1
Paid in Capital in Excess of Par- Common Stock    6,240,000
(To record issuance of 240,000 common shares of $1 par value @ $27 per share)
b Treasury Stock       151,200 =5,400*28
Cash       151,200
(To record purchase of 5,400 previously issued common shares @ $28 per share)